Estimate the number of steps you would have to take to walk a distance equal to the circumference of the Earth. (We estimate that the length of a step for an average person is about 18 inches, or roughly 0.5 m. The radius of the Earth is 6.38 ✕ 106 m.)
1
Expert's answer
2018-02-18T08:21:08-0500
Let the average step length is l=0.50 m.
The Earth radius is R=6.38×〖10〗^6 m.
The circumference of Earth L= 2πR= 2π(6.38 ×〖10〗^6)
So, number of steps is
N=L/l=2πR/l=(2π(6.38 ×〖10〗^6))/0.50=8.02×〖10〗^7